in a class of final year student in a school, 350 offer chemistry and 200 offer physics. if 90 student offer neither chemistry nor physics;


i: how many students are offering both subject.


ii: how many students are offering only one subject.


iii: how many students are offering at least one subject


1
Expert's answer
2021-11-04T17:17:58-0400

Out of 500 students in a final year of a senior secondary, 350 offer chemistry and 200 offer physics.

i. Let "x=" the number of students offering both subjects



Then


"500=x+350-x+200-x+90"

"x=140"

140 students are offering both subjects.


ii.


"350-140=210"

"200-140=60"

"210+60=270"

270 students are offering only one subject.


iii.


"500-90=410"

410 students are offering at least one subject.